OpenStudy (anonymous):

The area of a rectangle must be at least 135 square feet. If the length is 9 feet, find the minimum for the rectangle's width.

hero (hero):

A = lw 135 = 9w w = 135/9 w = 15

hero (hero):

\[w \ge 15\]